The doors into the clean room have been sealed by temporary walls for the renovation project which will start in the room next door. Below is a picture of the wall which has sheet rock on top of a plastic cover surrounding metal studs. The wall has been calked on the outside to prevent dust from coming through.
Also note that the HVAC temperature setpoint was changed to 68° F around noon.
2 - 480 V outlets have been installed.
Cleanroom fans were turned on this afternoon around 2:30 PM.
The chamber was wiped down and vacuumed.
All of the pre-tensioners have been set to a spring height of 1.75".
9/26/11 - Stringing Begins
Temperature fluctuations were within ± 2° during the stringing
HVAC temperature setting was dropped to 67° from 68° and the Δt was changed from ±2° to ±4° at around 1 PM.
Due to the late start this afternoon, we were only able to string 3 guard wires.
Temperature fluctuations were within ± 1° during the stringing
With 3 stringers rotating, we were able to string 45 Guard wires today.
A strip door was installed between the cleanroom area and the beam lab counting room. Several ducting pipes that are connected to adjacent rooms were also sealed off.
69 Guard wires were strung today.
The graph below was drawn at 6:32 pm on 9/29/11. It seems time stamp 0:00 corresponds to 6pm. It appears that the stringers enter the clean room around 8am (14:00) and warm up the clean room by almost 1 degree. TEMP4 is a thermometer on the vinyl wall right next to the chamber that is currently being strung. Temp1 is the middle of the clean room. Temp2 and Temp3 are on the opposite wall to the drift chamber currently being strung.
Since the previous change of our Δt to ±4° only changed the amplitude of our temperature change and didn't change our frequency, our Δt was changed back to ±2° at 2:30 PM.
91 guard wires were strung today.
The HVAC temperature set point was increased to 68° F at 4:15 PM.
The HVAC temperature set point was increased to 68° F yesterday at 4:15 PM.
The overnight temperatures are shown below (image taken at 8:45 am on 9/30/11)
Guard wires for SL2 were completed at 11:30 AM.
Pictures below show us rotating the chamber in preparation for stinging SL1 guard wires.
We started stringing guard wires for SL1. Total wires strung today: 51.
- Average Temperature for the First week of stringing